A family of 5 comes to the amusement park. There are two adults and three children (under the age of 12). They qualify for a special family rate of $174.45. The clerk tells them that a child's ticket always costs $11 less than an adult's ticket. How much is each adult ticket and how much is each child's ticket? (the original adult ticket w/o the family plan id $44.95) I cant figure out the variables and how to solve this.
Answer by mananth(16949)

Family
2 adults
3 kids
Family rate 174.45
adult ticket = 44.95
$ 11 less always
Child's ticket = 33.95
adult ticket cost x
child ticket cost y
y==33.95
2x+3y=174.45
y=33.95
2x+101.85=174.95
2x=174.95-101.85
2x=73.10
x=36.55 adults ticket cost
m.ananth@hotmail.ca